An empty Erlenmeyer flask weighs 256.1 g. When filled with water (d = 1.00 g/cm3), the flask and its contents weigh 438.3 g.
1
Expert's answer
2020-08-26T03:08:06-0400

∆m = 438.3 g – 256.1 g = 182.2 g

"d = \frac{m}{V}"

"V = \frac{m}{d}"

"V = \frac{182.2 g}{1\frac{g}{cm^3}} = 182.2 cm^3"

Answer: volume of the Erlenmeyer flask is 182.2 ml